OpenStudy (anonymous):

..The time t required to drive a certain distance varies inversely with the speed r. If it takes 2 hours to drive the distance at 30 miles per hour, how long will it take to drive the same distance at 50 miles per hour? A. 60 hours B. 1.2 hours C. 750 hours D. about 3.33 hours

OpenStudy (anonymous):

The time required to drive any distance would vary inversely with the (presumably constant) speed. You don't even need to compute this. Driving two hours at 30 mph will get you how far? Is this a roughly the distance you'd go at 50 mph for an hour? Which answer is close?

Directrix (directrix):

B. 1.2 hours t = d/r 2 = d/30 d = 60 miles ------- t = 60/50 t = 6/5 or 1.2 hours --> Option B